# Probing forced responses and causality in data-driven climate emulators: conceptual limitations and the role of reduced-order models

### v5

@@ IV.2.3 Interpretability and causal understanding through ensemble mean responses
The sensitivity map $\mathbf{S}$ is shown in Figure 7. The obtained sensitivity map $\mathbf{S}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [FDTPatternEffect]), with other data-driven approaches (see Figure 7(b) in [kang2023recent] and Figure 2 in [senneNN]), and with the one obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [blochjohnsonetal2024], see Figure 5(a) in [Dong2019AttributingPacific]). In the latter case, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments. For additional statistical reconstruction approaches to the pattern effect that yield structures similar to those in Figure 7, see the recent analysis by Fredericks et al. (2025) [Leif].
@@ V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
These results highlight intrinsic limitations of purely data-driven modeling of high-dimensional, multiscale climate systems, where the full set of relevant variables is unobserved or even unknown  [^2] A brief remark: in partially observed settings, one might be tempted to appeal to Takens’ embedding theorem [Takens]. This is not a practical option: the theorem does not extend to stochastic systems, and it becomes infeasible in high-dimensional settings [Cecconi, Baldovin, Lucente].. This is especially true for emulators built to capture real-world dynamics from observational data, where observations are inherently partial. In this context, the “correct” choice of variables, scales, and stochastic parameterizations depends ultimately on the specific scientific goals [Smith2001]. This perspective helps explain the continued success of Linear Inverse Models (LIMs) proposed by Penland, C. (1989) [Penland89], which rely on careful selection of climate fields and preprocessing choices to capture targeted aspects of climate variability; see, for example, [Penland95, PENLAND1996534, Penland2025] among others.
@@ VI Conclusion
Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system. This raises the fundamental challenge of identifying the proper variables (both fields and spatiotemporal scales), a fundamental problem in physics emphasized in classical work [Onsager] and revisited in recent studies [Lucente]. Such choices are not unique, especially in the case of observational data, and are largely determined by the study’s scientific goal rather than by the pursuit of an unattainable perfect model [Smith2001]. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is dominated by a few recurrent spatiotemporal patterns, or modes of variability, and their interactions. Thus, at least for certain tasks, the multiscale structure of the climate system can be an advantage, allowing us to focus on the coarse-grained dynamics while treating unresolved, fast processes stochastically. Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play. Such data-driven modeling approach has a long history in climate science [LucariniChekroun], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [Penland89, Penland95, KONDRAENSO] rather than forced responses. A promising perspective is to build on such established reduced-order approaches, including those developed by Majda and Qi [majdaSIAM, QiMajdaChaos], while leveraging flexible neural networks architectures to discover low-dimensional latent representations [latentPierre, PierreLatent] and fit deterministic and stochastic components. Furthermore, tools from response theory such as the FDT [FabCausal, LudovicoFabriAndre, ValerioData2] can be used to guide model development and eventually lead to reduced-order models that efficiently capture responses in the mean and variance to external perturbations, advancing causal inference and attribution analyses in climate science.
@@ A.2 Causal interpretation of the response operator
Consider a $N$ dimensional dynamical system of the kind:
$\begin{split}\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{N});~\text{with}~i=1,2,...,N.\end{split}$ (14)
For stochastic systems, causal links are contained in the deterministic drift $f_{i}$, so the discussion below applies to both deterministic and stochastic dynamics. In this setting, the existence of a direct causal link $x_{j}\rightarrow x_{k}$ is determined by whether $x_{j}$ appears in the arguments of $f_{k}$. More formally, the absence of a direct causal link $x_{j}\rightarrow x_{k}$ is expressed by:
$\frac{\partial f_{k}}{\partial x_{j}}=0.$
As a concrete example, we return to the triad system in Eq. (6), for which the underlying direct causal links are known exactly. This example provides a straightforward, intuitive illustration of how perturbation experiments reveal causal interactions. Figure 8 reports both (a) its functional form and (b) its causal graph, which encode the same information.
@@
  • Any perturbation applied to $x_{2}$ will affect $x_{2}$ itself and $x_{1}$ after one time step, but not $x_{3}$. Hence, we infer $\partial f_{3}/\partial x_{2}=0$.
  • A perturbation applied to $x_{3}$ will affect $x_{3}$ and $x_{1}$, but not $x_{2}$, so $\partial f_{2}/\partial x_{3}=0$.
@@ Acknowledgements.
Most of the ideas discussed in this work stem from collaborations and exchanges with colleagues and friends. I am grateful for the supportive and engaging environment of the Zanna Group at Courant, as well as for interactions within the broader M2Lines project. I thank Laure Zanna and Carlos Fernandez-Granda for their feedback during early stages of the project. I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system. I acknowledge an insightful email exchange with Prof. Tony Roberts, who drew my attention to recent advances on stochastic model reduction. I thank two anonymous reviewers for their very useful and constructive comments. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.

### v6

@@ IV.2.3 Interpretability and causal understanding through ensemble mean responses
The sensitivity map $\mathbf{S}$ is shown in Figure 7. The obtained sensitivity map $\mathbf{S}$ is in qualitative agreement with previous results in response theory (Figure 2(a) in [FDTPatternEffect]), with other data-driven approaches (see Figure 7(b) in [kang2023recent] and Figure 2 in [senneNN]), and with the one obtained by directly perturbing the SST boundary condition of atmosphere-only climate models (see third row of Figure 2 in [blochjohnsonetal2024], see Figure 5(a) in [Dong2019AttributingPacific]). In the case of atmosphere-only climate models, the maps exhibit fewer spatial features due to the coarse resolution of the perturbation patches (spanning tens of degrees) in numerical models experiments. Focusing on smaller perturbation patches (approaching grid-scale perturbations) yields more finely resolved spatial features in climate-model experiments (see Figure 3 in [Leif]), a point sometimes invoked to motivate higher-resolution modeling for forced responses. Our results, however, show that similarly detailed sensitivity patterns can be obtained with a reduced-order model using only 20 principal components. This suggests that high spatial resolution is not required for this problem; what matters is representing the system in an appropriate reduced (coarse-grained) representation. For additional statistical reconstruction approaches to the pattern effect that yield structures similar to those in Figure 7, see the recent analysis by Fredericks et al. (2025) [Leif].
@@ V.1 Limits of data-driven-only climate emulators and the role of stochastic approaches
These results highlight intrinsic limitations of purely data-driven modeling of high-dimensional, multiscale climate systems, where the full set of relevant variables is unobserved or even unknown. This is especially true for emulators built to capture real-world dynamics from observational data, where observations are inherently partial. In this context, the “correct” choice of variables, scales, and stochastic parameterizations depends ultimately on the specific scientific goals [Smith2001]. This perspective helps explain the continued success of Linear Inverse Models (LIMs) proposed by Penland, C. (1989) [Penland89], which rely on careful selection of climate fields and preprocessing choices to capture targeted aspects of climate variability; see, for example, [Penland95, PENLAND1996534, Penland2025] among others.
@@ VI Conclusion
Modeling the climate system from data alone is constrained by observing only a subset of variables of a multiscale, high-dimensional dynamical system. This raises the fundamental challenge of identifying the proper variables (both fields and spatiotemporal scales), a fundamental problem in physics emphasized in classical work [Onsager] and revisited in recent studies [Lucente]. Such choices are inherently non-unique, especially because the climate system lacks clear space- and time-scale separation, and are therefore largely determined by the study’s scientific goal rather than by the pursuit of an unattainable perfect model [Smith2001]. In contrast to numerical weather prediction, which demands high spatial and temporal resolution, low-frequency climate variability is dominated by a few recurrent spatiotemporal patterns, or modes of variability, and their interactions. Thus, at least for certain tasks, the multiscale structure of the climate system can be an advantage, allowing us to focus on the coarse-grained dynamics while treating unresolved, fast processes stochastically. Traditional reduced-order models stand out from general-purpose emulators by targeting specific climate problems from the outset, guided by prior literature, physical insight, and a clear understanding of the relevant variables and scales at play. Such data-driven modeling approach has a long history in climate science [LucariniChekroun], though most efforts have emphasized reproducing stationary statistics or short-term forecasting [Penland89, Penland95, KONDRAENSO] rather than forced responses. A promising perspective is to build on such established reduced-order approaches, including those developed by Majda and Qi [majdaSIAM, QiMajdaChaos], while leveraging flexible neural networks architectures to discover low-dimensional latent representations [latentPierre, PierreLatent] and fit deterministic and stochastic components. Furthermore, tools from response theory such as the FDT [FabCausal, LudovicoFabriAndre, ValerioData2] can be used to guide model development and eventually lead to reduced-order models that efficiently capture responses in the mean and variance to external perturbations, advancing causal inference and attribution analyses in climate science.
@@ A.2 Causal interpretation of the response operator
Consider a high-dimensional, stochastic dynamical system of the kind:
$\begin{split}\frac{dx_{i}}{dt}=f_{i}(x_{1},x_{2},...,x_{N})+\text{stochastic forcing};\\ ~\text{with}~i=1,2,...,N~.\end{split}$ (14)
In this setting, the existence of a direct causal link $x_{j}\rightarrow x_{k}$ is determined by whether $x_{j}$ appears in the arguments of $f_{k}$. More formally, the absence of a direct causal link $x_{j}\rightarrow x_{k},j\neq k$ is expressed by:
$\langle\frac{\partial f_{k}(\mathbf{x})}{\partial x_{j}}\rangle=0.$
We adopt the view that causal estimates in real-world, high-dimensional stochastic systems are most naturally, and most practically, formulated as ensemble averages, making the ensemble-averaged Jacobian the appropriate structural object for direct causal links. As a concrete example, we return to the triad system in Eq. (6), for which the underlying direct causal links are known exactly. This example provides a straightforward, intuitive illustration of how perturbation experiments reveal causal interactions. Figure 8 reports both (a) its functional form and (b) its causal graph, which encode the same information.
@@
  • Any perturbation applied to $x_{2}$ will affect, on average, $x_{2}$ itself and $x_{1}$ after one time step, but not $x_{3}$. Hence, we infer $\langle\partial f_{3}(\mathbf{x})/\partial x_{2}\rangle=0$.
  • A perturbation applied to $x_{3}$ will affect, on average, $x_{3}$ and $x_{1}$, but not $x_{2}$, so $\langle\partial f_{2}(\mathbf{x})/\partial x_{3}\rangle=0$.
@@ Acknowledgements.
Most of the ideas discussed in this work stem from collaborations and exchanges with colleagues and friends. I am grateful for the supportive and engaging environment of the Zanna Group at Courant, as well as for interactions within the broader M2Lines project. I thank Laure Zanna and Carlos Fernandez-Granda for their feedback during early stages of the project. I am especially grateful to Pasha Perezhogin, Rory Basinski, Chris Pedersen, and Andre Souza for their thoughtful feedback and for many enriching discussions on data-driven modeling of the Climate system. I acknowledge an insightful email exchange with Prof. Tony Roberts, who drew my attention to recent advances on stochastic model reduction. I thank two anonymous reviewers for their very useful and constructive comments. This research received support through Schmidt Sciences, LLC, under the M2LInES project. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.